# The $\Lambda_{b}\to\Lambda$ transition form factors in perturbative QCD approach

### v1

@@ III Light-cone distribution amplitudes
$a_{0}^{f_{+}}$ $\phantom{-}0.419\pm 0.031$ $a_{0}^{g_{\perp}}$ $0.358\pm 0.027$
$a_{1}^{f_{+}}$ $1.04\pm 0.142$ $a_{1}^{g_{\perp}}$ $0.890\pm 0.123$
$a_{0}^{f_{0}}$ $\phantom{-}0.371\pm 0.032$ $a_{0}^{h_{+}}$ $\phantom{-}0.490\pm 0.050$
$a_{1}^{f_{0}}$ $0.830\pm 0.146$ $a_{1}^{h_{+}}$ $1.150\pm 0.208$
$a_{0}^{f_{\perp}}$ $\phantom{-}0.520\pm 0.040$ $a_{0}^{h_{\perp}}$ $\phantom{-}0.385\pm 0.034$
$a_{1}^{f_{\perp}}$ $1.290\pm 0.167$ $a_{1}^{h_{\perp}}$ $0.890\pm 0.156$
$a_{0}^{g_{+}}$ $\phantom{-}0.356\pm 0.023$ $a_{0}^{\widetilde{h}_{+}}$ $\phantom{-}0.339\pm 0.029$
$a_{1}^{g_{+}}$ $0.760\pm 0.107$ $a_{1}^{\widetilde{h}_{+}}$ $0.810\pm 0.134$
$a_{0}^{g_{0}}$ $\phantom{-}0.397\pm 0.029$ $a_{0}^{\widetilde{h}_{\perp}}$ $0.343\pm 0.029$
$a_{1}^{g_{0}}$ $\phantom{-}-0.930\pm 0.137$ $a_{1}^{\widetilde{h}_{\perp}}$ $0.700\pm 0.130$
Table 7: Central values and uncertainties of the nominal form factor parameters.
Figure 3: Form factors of the entire kinematic range, and the bands show the uncertainty.
@@
$\frac{\mathrm{d}^{4}\Gamma}{\mathrm{d}q^{2}\>\mathrm{d}\cos\theta_{\ell}\>\mathrm{d}\cos\theta_{\Lambda}\>\mathrm{d}\phi}$ $=$ $\frac{3}{8\pi}\bigg{[}\hskip 12.05553pt\big{(}K_{1ss}\sin^{2}\theta_{\ell}+\,K_{1cc}\cos^{2}\theta_{\ell}+K_{1c}\cos\theta_{\ell}\big{)}$ (38)
@@
where the angles $\theta_{\Lambda}$ and $\theta_{\ell}$ describe the polarization directions of the proton and negatively charged lepton, respectively. $K_{n\lambda}$ with $n=1,\dots,4$ and $\lambda=s,ss,c,cc,sc$ are functions of $q^{2}$, and its explicit expressions can be found in Eqs.(3.29)-(3.32) of Ref. Boer:2014kda . $\phi$ is the azimuthal angle between the decay planes of $p\pi$ and $\ell\ell$. By integrating over the three angles appearing in Eq. (38), we can obtain the $q^{2}$-differential decay rate as
@@
In addition to the differential decay rate, we also define the normalized angular observables.
$\hat{K}_{i}=\frac{K_{i}}{\mathrm{d}\Gamma/\mathrm{d}q^{2}},$ (40)
@@
Figures 4 and 5 present our theoretical predictions for, respectively, the differential branching fraction of the $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$ decay and the angular observables in the quasi four-body decay $\Lambda_{b}\to\Lambda(\to p\pi^{-})\mu^{+}\mu^{-}$. In the current analysis, we focus exclusively on the dominant theoretical uncertainty arising from the form factors, while other potential sources of uncertainty are neglected as they are expected to be subdominant in comparison. By integrating the differential branching fraction over the kinematic range $(2m_{\mu})^{2}\leq q^{2}\leq(M_{\Lambda_{b}}-m_{\Lambda})^{2}$, we obtain the total branching fraction for the $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$ decay:
$\int_{(2m_{\mu})^{2}}^{(M_{\Lambda_{b}}-m_{\Lambda})^{2}}\frac{\mathrm{d}\mathcal{B}}{\mathrm{d}q^{2}}\mathrm{d}q^{2}=\left(1.24^{+0.25}_{-0.22}\right)\times 10^{-6}.$ (45)
where the integration is performed numerically across the full physical phase space. This prediction agrees well with the current experimental measurement of $(1.08\pm 0.28)\times 10^{-6}$ reported by the PDG ParticleDataGroup:2024cfk .
As evident from Fig. 4, the theoretical uncertainty in our result stemming from the form factor calculations is large. Future improvements in determining the light-cone distribution amplitudes (LCDAs) of both the $\Lambda_{b}$ and $\Lambda$ baryons could significantly reduce this uncertainty. The angular observables presented in Fig. 5, being defined as ratios of decay rates, exhibit less sensitivity to these uncertainties due to partial cancellations between numerator and denominator.
Figure 4: $\Lambda_{b}\to\Lambda\>\mu^{+}\mu^{-}$ differential branching fraction, compared to the experimental data from LHCb LHCb:2015tgy .
Figure 5: $\Lambda_{b}\to\Lambda(\to p^{+}\pi^{-})\mu^{+}\mu^{-}$ angular observables, compared to the experimental data from LHCb LHCb:2015tgy .
@@ VI Summary
In this paper, we calculate the $\Lambda_{b}\to\Lambda$ transition form factors by the PQCD approach, and our results for the transition form factors in $q^{2}$ region are close to those in other theoretical approaches within errors. In our calculation, we consider the contributions of $\Lambda_{b}$ LCDAs up to twist-4 and $\Lambda$ LCDAs up to twist-6, We find that the combination term of $\Lambda_{b}$ twist-4 and $\Lambda$ twist-5 provides the largest contribution to the form factors, this indicates that the contributions of higher-twist terms in the LCDAs of baryons should be properly taken into account in the study of $\Lambda_{b}$ baryon decays. In addition, we obtained the form factors over the full kinematic region by performing a global fit that combines the PQCD results in the low-$q^{2}$ region with the lattice QCD results in the high-$q^{2}$ region. Using these form factors, we calculated the differential branching fraction and angular observables. Our results are consistent with most theoretical predictions and experimental data within the uncertainty range. Furthermore, due to the adoption of the latest $\Lambda_{b}\to p\pi$ asymmetry parameter, our prediction for hadron-side asymmetry is in better agreement with experimental data.
@@ Acknowledgment
The authors would like to thank Dr. Dong-Hao Li for helpful discussions and valuable suggestions. The work is supported by the National Natural Science Foundation of China (Grant No. 12275067 and 12335003), Science and Technology R&D Program Joint Fund Project of Henan Province (Grant No. 225200810030), Science and Technology Innovation Leading Talent Support Program of Henan Province, National Key R&D Program of China (Grant No. 2023YFA1606000), and the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2023-stlt01, lzujbky-2024-oy02 and lzujbky-2025-eyt01).
Appendix A:Factorization formulas

### v2

@@ III Light-cone distribution amplitudes
$a_{0}^{f_{+}}$ $\phantom{-}0.418\pm 0.031$ $a_{0}^{g_{\perp}}$ $0.357\pm 0.027$
$a_{1}^{f_{+}}$ $1.21\pm 0.15$ $a_{1}^{g_{\perp}}$ $0.99\pm 0.14$
$a_{0}^{f_{0}}$ $\phantom{-}0.370\pm 0.032$ $a_{0}^{h_{+}}$ $\phantom{-}0.49\pm 0.05$
$a_{1}^{f_{0}}$ $1.00\pm 0.16$ $a_{1}^{h_{+}}$ $1.37\pm 0.23$
$a_{0}^{f_{\perp}}$ $\phantom{-}0.52\pm 0.04$ $a_{0}^{h_{\perp}}$ $\phantom{-}0.386\pm 0.034$
$a_{1}^{f_{\perp}}$ $1.50\pm 0.19$ $a_{1}^{h_{\perp}}$ $1.04\pm 0.17$
$a_{0}^{g_{+}}$ $\phantom{-}0.356\pm 0.023$ $a_{0}^{\widetilde{h}_{+}}$ $\phantom{-}0.338\pm 0.029$
$a_{1}^{g_{+}}$ $0.89\pm 0.13$ $a_{1}^{\widetilde{h}_{+}}$ $0.91\pm 0.15$
$a_{0}^{g_{0}}$ $\phantom{-}0.397\pm 0.029$ $a_{0}^{\widetilde{h}_{\perp}}$ $0.342\pm 0.029$
$a_{1}^{g_{0}}$ $\phantom{-}-1.07\pm 0.15$ $a_{1}^{\widetilde{h}_{\perp}}$ $0.85\pm 0.15$
Table 7: Central values and uncertainties of the fit parameters.
Figure 3: Form factors at the full kinematic range.
@@
$\frac{\mathrm{d}^{4}\Gamma}{\mathrm{d}q^{2}\>\mathrm{d}\cos\theta_{\ell}\>\mathrm{d}\cos\theta_{\Lambda}\>\mathrm{d}\phi}$ $=$ $\frac{3}{8\pi}\bigg[\hskip 10.84996pt\big(K_{1ss}\sin^{2}\theta_{\ell}+\,K_{1cc}\cos^{2}\theta_{\ell}+K_{1c}\cos\theta_{\ell}\big)$ (35)
@@
where the angles $\theta_{\Lambda}$ and $\theta_{\ell}$ describe the polarization directions of the proton and negatively charged lepton, respectively. $K_{n\lambda}$ with $n=1,\dots,4$ and $\lambda=s,ss,c,cc,sc$ are functions of $q^{2}$, and its explicit expressions can be found in Eqs.(3.29)-(3.32) in Ref. Boer:2014kda . $\phi$ is the azimuthal angle between the decay planes of $p\pi$ and $\ell^{+}\ell^{-}$. By integrating over the three angles appearing in Eq. (35), we can obtain the $q^{2}$-differential decay rate as,
@@
In addition to the differential decay rate, we also define the normalized angular observables,
$\hat{K}_{i}=\frac{K_{i}}{\mathrm{d}\Gamma/\mathrm{d}q^{2}}.$ (37)
@@
Differential branching ratio of the $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$ decay obtained in this work is shown in Fig. 4. Fig. 5 presents the angular observables in the quasi four-body decay $\Lambda_{b}\to\Lambda(\to p\pi^{-})\mu^{+}\mu^{-}$. In the current analysis, we focus only on the dominant theoretical uncertainties arising from the $\Lambda_{b}\to\Lambda$ form factors, the other potential sources of uncertainty are neglected as they are expected to be subdominant. By integrating the differential branching fraction over the kinematic range $(2m_{\mu})^{2}\leq q^{2}\leq(M_{\Lambda_{b}}-m_{\Lambda})^{2}$, we obtain the total branching fraction for the $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$ decay,
$\int_{(2m_{\mu})^{2}}^{(M_{\Lambda_{b}}-m_{\Lambda})^{2}}\frac{\mathrm{d}\mathcal{B}}{\mathrm{d}q^{2}}\mathrm{d}q^{2}=\left(1.08^{+0.15}_{-0.14}\right)\times 10^{-6}.$ (42)
The above prediction is in good agreement with the experimental measurement, $(1.08\pm 0.28)\times 10^{-6}$ ParticleDataGroup:2024cfk .
As shown in Fig. 4, the theoretical uncertainty for the differential branching ratio stemming from the form factors is large. Future improvements in determining the LCDAs of both $\Lambda_{b}$ and $\Lambda$ baryons could significantly reduce uncertainty of the branching ratio. The angular observables presented in Fig. 5 exhibit less sensitivity to uncertainties of the form factors due to partial cancellation in their definitions.
Figure 4: Differential branching fraction of the $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$ decay. The cyan curve and the lighter band correspond to the central values and theoretical uncertainties obtained in this work, respectively, compared with the experimental data from LHCb LHCb:2015tgy .
Figure 5: Angular observables of the $\Lambda_{b}\to\Lambda(\to p^{+}\pi^{-})\mu^{+}\mu^{-}$ decay obtained in this work, where the bands represent the corresponding theoretical uncertainties, compared with the experimental data from LHCb LHCb:2015tgy .
@@ VI Summary
In this paper, we calculate the $\Lambda_{b}\to\Lambda$ transition form factors in the PQCD approach. The obtained form factors are close to those from other theoretical approaches within uncertainties. In our calculation, we consider the contributions of the $\Lambda_{b}$ LCDAs up to twist-4 and $\Lambda$ LCDAs up to twist-6, and find that high-power distribution amplitudes provide the dominant contributions to the form factors, indicating that the high-power corrections should be properly taken into account in the studies of baryon decays. In addition, we obtained the form factors at the full kinematic region by performing a z-series extrapolation, in which we combine the PQCD predictions at the low-$q^{2}$ region and the lattice QCD results at the $q^{2}_{max}$. We then evaluate the differential branching fraction of the $\Lambda_{b}\to\Lambda\mu^{+}\mu^{-}$ decay and angular observables of the quasi four-body $\Lambda_{b}\to\Lambda(\to p\pi^{-})\mu^{+}\mu^{-}$ decay. Our results are consistent with the most theoretical predictions and the experimental measurements within uncertainties.
@@ Acknowledgment
The authors would like to thank Dr. Dong-Hao Li for helpful discussions and valuable suggestions. The work is supported by the National Natural Science Foundation of China (Grant No. 12275067, 12335003 and 12505113), Science and Technology R&D Program Joint Fund Project of Henan Province (Grant No. 225200810030), Science and Technology Innovation Leading Talent Support Program of Henan Province, National Key R&D Program of China (Grant No. 2023YFA1606000), and the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2023-stlt01, lzujbky-2024-oy02 and lzujbky-2025-eyt01).
Appendix: Factorization formulas